# Homeomorphisms of continua through projective Fraïssé limits

### v1

@@ 5. A homeomorphism of the pseudoarc that is not conjugate to an automorphism of the pre-pseudoarc
In this section will concern $\mathcal{P}$}, the class of finite connected linear graphs with all edge-preserving surjective homomorphisms (note that if a graph-homomorphism between two finite linear graphs is onto it is automatically surjective on the edges).
@@
Let $P$ be the pseudoarc, which we identify with $\rm pr$$\mathbb{P}$}], the natural projection of the projective Fraïssé limit of the class of finite linear graphs. There exists a homeomorphism $\sigma\in{\rm Homeo}(P)$ that is not conjugate to any $\tau\in{\rm Aut}({\mathbb{P}})$.
The following is a key lemma; it provides a sufficient condition for a homeomorphism to not be conjugate to any automorphism of $\mathbb{P}$}. Our argument hinges on the rigidity of automorphisms, that they preserve the algebra of regular open sets associated with the projections onto finite linear graphs. Our example that will be constructed in Proposition 5.4 will satisfy the premises.
@@
Suppose that $\sigma:P\to P$ is a homeomorphism such that there exists a subcontinuum $C\subseteq P$, $x\neq y\in C$ with the property that
  1. (1)
    $\sigma{\restriction}C={\rm id}_{C}$, and
  2. (2)
    whenever $U_{1},U_{2},U_{3}$ are pairwise disjoint regular open sets in $P$ with
    1. (a)
      $\bigcup_{i=1}^{3}{\rm cl}(U_{i})=P$,
    2. (b)
      $\bigcap_{i=1}^{3}{\rm cl}(U_{i})=\emptyset$ (that is, each point has a neighborhood which intersects only two of the $U_{i}$’s),
    3. (c)
      $C\subseteq{\rm int}({\rm cl}(V_{2})\cup{\rm cl}(V_{3}))$ (equivalently, ${\rm cl}(U_{1})\cap C=\emptyset$, $C$ is positive distance from $U_{1}$),
    4. (d)
      $x\in U_{2}$, $y\in U_{3}$,
    then $\sigma(U_{2})\cap U_{3}\neq\emptyset$, or $\sigma(U_{3})\cap U_{2}\neq\emptyset$.
Then $\sigma$ is not conjugate to any member of Aut(($\mathbb{P}$}).
@@
Note that the above property of $\sigma$ is invariant under conjugation by a homeomorphism, therefore it suffices to show that representing $P$ as a quotient of the prespace $\mathbb{P}$}, necessarily $\sigma\notinAut$$({\mathbb{P}}$).
[…]
Suppose that $\sigma$ is an automorphism, which formally means that there exists some $\tilde{\sigma}$$\mathbb{P}$$\to{\mathbb{P}}$ with ${\rm pr}(\tilde{\sigma})=\sigma$ (where ${\rm pr}:{\rm Aut}{\mathbb{P}}\to{\rm Homeo}(P)$ is the canonical embedding; in other words, the equality $\sigma\circ{\rm pr}={\rm pr}\circ\tilde{\sigma}$ holds).
[…]
For each $j$ we let $X_{j}=\pi_{j}\circ{\rm pr}^{-1}(x)$, $Y_{j}=\pi_{j}\circ{\rm pr}^{-1}(y)$, so $X_{j},Y_{j}\subseteq I_{j}$, $1\leq|X_{j}|,|Y_{j}|\leq 2$. Let $j_{0}$ be large enough so that $X_{j_{0}}\cup Y_{j_{0}}$ spans a not connected subgraph in $J_{j_{0}}\subseteq I_{j_{0}}$ (i.e. there is a gap in between), moreover, each connected component of $I_{k}\setminus(X_{k}\cup Y_{k})$ has at least two elements.
Pick a large enough $k$ so that $\pi_{j_{0}}\circ\tilde{\sigma}$ factors through $\pi_{k}$, i.e. for some epimorphism $\alpha:I_{k}\to I_{j_{0}}$ we have
(5.1) $\pi_{j_{0}}\circ\tilde{\sigma}=\alpha\circ\pi_{k}.$
Claim 5.3.
$\alpha{\restriction}J_{k}=\pi_{j_{0},k}$.
First we argue that the claim will finish the proof of the lemma. Pick disjoint subsets $X^{\prime},Y^{\prime}\subseteq J_{j_{0}}$ with $X^{\prime}\cup Y^{\prime}=J_{j_{0}}$, $X_{j_{0}}\subseteq X^{\prime}$ and $Y_{j_{0}}\subseteq Y^{\prime}$. Let $X^{+}=\pi_{j_{0},k}^{-1}(X^{\prime})\cap J_{k}$, $Y^{+}=\pi_{j_{0},k}^{-1}(Y^{\prime})\cap J_{k}$, in particular, $X^{+}\cup Y^{+}=J_{k}$, and
(5.2) $\pi_{j_{0},k}[X^{+}]\cap\pi_{j_{0},k}[Y^{+}]=\emptyset,$
(5.3) $X_{k}\subseteq X^{+},\ \ Y_{k}\subseteq Y^{+},$
Let $V_{1}=I_{k}\setminus J_{k}$ be the union of the two complementer intervals. Now $W_{1}={\rm pr}(\pi_{k}^{-1}(V_{1})),W_{2}={\rm pr}(\pi_{k}^{-1}(X^{+})),W_{3}={\rm pr}(\pi_{k}^{-1}(Y^{+}))$ are closed sets that cover $P$. The fact that $X^{+}\cup Y^{+}=J_{k}=\pi_{k}[{\rm pr}^{-1}(C)]$ together with $V_{1}\cap J_{k}=\emptyset$ imply $C\cap W_{1}=\emptyset$. Also, it follows from (5.3) and the way $X_{k},Y_{k}$ were defined that
$x\notin W_{1}\cup W_{3},$
$y\notin W_{1}\cup W_{2}.$
Recalling Lemma 2.1 and using that ${\mathcal{K}}={\mathcal{P}}$ is the class of finite linear graphs it is routine to check that for any $S\subseteq I_{k}$, letting $D_{S}\subseteq{\mathbb{P}}$ denote the points that project to $S$ (i.e. $D_{S}=\pi_{k}^{-1}(S)$),
${\rm int}({\rm pr}(D_{S}))={\rm pr}(D_{S})\setminus{\rm pr}(D_{I_{k}\setminus S}),$
and ${\rm cl}({\rm int}({\rm pr}(D_{S})))={\rm pr}(D_{S}))$, so $U_{i}={\rm int}(W_{i})$, $i=1,2,3$ are pairwise disjoint regular open sets satisfying the requirements above.
@@
We assume that $\sigma(U_{2})\cap U_{3}\neq\emptyset$ as the other case can be dealt with the same way. So pick $z\in{\mathbb{P}}$ witnessing this, i.e. ${\rm pr}(z)\in U_{2}$, $\sigma({\rm pr}(z))={\rm pr}(\tilde{\sigma}(z))\in U_{3}$. By the definition of $U_{2}$, we have ${\rm pr}(z)\in W_{2}\setminus(W_{1}\cup W_{3})$, so $\pi_{k}(z)\in X^{+}$ necessarily. Similarly, $\pi_{k}(\tilde{\sigma}(z))\in Y^{+}$, so
(5.4) $\pi_{j_{0}}(\tilde{\sigma}(z))=\pi_{j_{0},k}(\pi_{k}(\tilde{\sigma}(z)))\in\pi_{j_{0},k}[Y^{+}].$
But by (5.1) $\pi_{j_{0}}(\tilde{\sigma}(z))=\alpha(\pi_{k}(z))$, so by Claim 5.3
(5.5) $\pi_{j_{0}}(\tilde{\sigma}(z))=\alpha(\pi_{k}(z))=\pi_{j_{0},k}(\pi_{k}(z))\in\pi_{j_{0},k}[X^{+}].$
Finally, note that (5.4), (5.5) contradict (5.2).
Proof..
(Claim 5.3) If $u\in J_{k}$, $v^{\prime}=\alpha(u)\neq\pi_{j_{0},k}(u)$, then letting $u^{\prime}$ denote $\pi_{j_{0},k}(u)$, clearly $\pi_{j_{0},k}^{-1}(u^{\prime})\cap\pi_{j_{0},k}^{-1}(v^{\prime})=\emptyset$. We now claim the following.
[…]
    There exists $k^{+}\geq k$, such that one of the following two possibilities holds:
    • for some $u^{+}\in J_{k^{+}}$ we have $\pi_{k,k^{+}}(u^{+})=u$, but $\neg(u^{+}R\pi_{j_{0},k^{+}}^{-1}(v^{\prime}))$,
    • for some $v^{+}\in J_{k^{+}}$ with $v=\pi_{k,k^{+}}(v^{+})$ we have $\pi_{j_{0},k}(v)=v^{\prime}$, $\alpha(v)=u^{\prime}$ and $\neg(v^{+}R\pi_{j_{0},k}^{-1}(u^{\prime}))$.
[…]
(5.6) $\neg(uR\pi_{j_{0},k}^{-1}(v^{\prime})),$
[…]
We assume that the first alternative fails. Then a standard argument shows that for some $(u_{i})_{i}$ with $u_{j_{0}}=u^{\prime}$, and $(v_{i})_{i}$ with $v_{j_{0}}=v^{\prime}$ we have $\tilde{\sigma}((u_{i})_{i})=(v_{i})_{i}$ and $u_{i}Rv_{i}$ for all $i$. Then (using that every node is of degree at most $1$ in the graph $R^{\mathbb{P}}$), $\tilde{\sigma}((v_{i})_{i})=(u_{i})_{i}$, in particular, $\alpha(v_{k})=u_{j_{0}}$.
[…]
@@
The only thing en route to a homeomorphism that is not conjugate to any element of Aut(($\mathbb{P}$}) is to construct one that satisfies the premise of Lemma 5.2.
@@
There exists an $\sigma\in{\rm Homeo}(P)$ as in Lemma 5.2.
@@
  • define a generic sequence $J_{0},J_{1},\ldots,J_{n},\ldots$ of finite linear graphs by hand (together with the bonding maps $\pi_{i,j}$, $i\leq j<$),

### v2

@@ 1. $(enumi)_{1}$
4(B). Density in ${\rm Aut}({\mathbb{P}},\bullet)^{\mathbb{N}}$
Consider the class $\mathcal{P}$$_{\bullet}$ of structures in the language consisting of one binary symbol $R$ and a constant symbol $c$. The structures of $\mathcal{P}$$_{\bullet}$ are finite sets with $R$ interpreted as a symmetric reflexive linear binary relation and the constant symbol $\bullet$ interpreted as one of the endpoints of the structure. So elements of $\mathcal{P}$$_{\bullet}$ are isomorphic to structures of the form
[…]
where $n\in{\mathbb{N}}$. Note that the enumeration of $I$ is not part of the structure. Morphisms in $\mathcal{P}$$_{\bullet}$ are all epimorphisms among structures in $\mathcal{P}$$_{\bullet}$. It follows from Lemma B.1 that the family $\mathcal{P}$$_{\bullet}$ forms a projective Fraïssé class. Using the argument from [9] that $\mathcal{P}$} is transitive, one checks that $\mathcal{P}$$_{\bullet}$ is transitive as well. Let ${\mathbb{P}}_{\bullet}\in({\mathcal{P}}_{\bullet})^{*}$ be its projective limit. We denote the underlying space of it by $\mathbb{P}$}. On this space we have interpretations $R^{\mathbb{P}}$ and $\bullet^{\mathbb{P}}$ of $R$ and $\bullet$, respectively. Then it is not difficult to check that the reduct $({\mathbb{P}},R^{\mathbb{P}})$ is the pre-pseudoarc, that is, it is isomorphic to the Fraïssé limit of $\mathcal{P}$} (in the category $\mathcal{P}$$^{*}$ derived from $\mathcal{P}$}). Therefore, the canonical quotient ${\mathbb{P}}/R^{{\mathbb{P}}}$ is homeomorphic to the pseudoarc and $\bullet^{P}=\bullet^{\mathbb{P}}/R^{\mathbb{P}}$ is a point in the pseudoarc $P$. Furthermore, the argument from [9] showing that ${\rm Aut}({\mathbb{P}})$ is dense in ${\rm Homeo}(P)$, gives that ${\rm Aut}({\mathbb{P}}_{\bullet})$ is dense in the set $\{h\in{\rm Homeo}(P)\mid h(\bullet_{P})=\bullet_{P}\}$.
Theorem 4.4.
  1. (i)
    For every $n\in\mathbb{N}$, the set of all $\bar{\gamma}\in{\rm Aut}({\mathbb{P}}_{\bullet})^{n}$ such that the orbit of $\bar{\gamma}$ under the diagonal conjugacy action by ${\rm Aut}({\mathbb{P}}_{\bullet})$ is dense in ${\rm Aut}(P_{\bullet})^{n}$ is comeager in ${\rm Aut}({\mathbb{P}}_{\bullet})^{n}$.
  2. (ii)
    The set of all $\bar{\gamma}\in{\rm Aut}({\mathbb{P}}_{\bullet})^{\mathbb{N}}$ such that the orbit of $\bar{\gamma}$ under the diagonal conjugacy action by ${\rm Aut}({\mathbb{P}}_{\bullet})$ is dense in ${\rm Aut}(P_{\bullet})^{\mathbb{N}}$ is comeager in ${\rm Aut}({\mathbb{P}}_{\bullet})^{\mathbb{N}}$.
The theorem will follow from the lemma below.
Lemma 4.5.
Suppose that $A,B,A^{\prime},B^{\prime}\in{\mathcal{P}}_{\bullet}$, there exist $A^{+},B^{+}\in{\mathcal{P}}_{\bullet}$ and $\alpha\in{\rm Epi}(A^{+},A)$, $\alpha^{\prime}\in{\rm Epi}(A^{+},A^{\prime})$, $\beta\in{\rm Epi}(B^{+},B)$, and $\beta^{\prime}\in{\rm Epi}(B^{+},B^{\prime})$ with the following property:
[…]
Proof..
Let
$A=\{0,\dots,m\},\;A^{\prime}=\{0,\dots,m^{\prime}\},\;B=\{0,\dots,n\},\;B^{\prime}=\{0,\dots,n^{\prime}\},$
with the constant in all these structures interpreted as 0.
We define $A^{+}$ with $R^{A^{+}}$. Let
$\begin{split}A^{\prime}_{0}&=\{(0,-1,j)\mid j\in A^{\prime}\}\\ A^{\prime}_{i}&=\{(i-1,1,j)\mid j\in A^{\prime}\}\cup\{(i,-1,j)\mid j\in A^{\prime}\},\;\hbox{ for }0<i\in A.\end{split}$
and we define $R^{A^{+}}$ on these sets to be the smallest reflexive and symmetric relation with
$(i,\epsilon,j)\,R^{A^{+}}(i,\epsilon,j^{\prime})\Leftrightarrow|j-j^{\prime}|\leq 1,\;\hbox{ for }i\in A\hbox{ and }\epsilon\in\{-1,1\},$
and
$(i-1,1,0)\,R^{A^{+}}(i,-1,0),\;\hbox{ for }0<i\in A.$
Note that with these definitions $A_{0}^{\prime}$ is a reflexive linear graph with $m^{\prime}+1$ many elements and $A_{i}^{\prime}$, $0<i\in A$, is a reflexive linear graph with $2(m^{\prime}+1)$ many elements.
Let now $A^{+}$ be the union
$\bigcup_{i\in A}A^{\prime}_{i}$
with the point $(i-1,-1,m^{\prime})\in A^{\prime}_{i-1}$ identified with the point $(i-1,1,m^{\prime})\in A^{\prime}_{i}$ for each $0<i\in A$. Let $R^{A^{+}}$ be the relation on $A^{+}$ that is induced from $R^{A^{+}}$ defined on $A^{\prime}$ and $A^{\prime}_{i}$ for $i\in A$, $i<m$. We interpret the constant as $(0,-1,0)\in A^{\prime}_{0}$.
The definitions of $B^{+}$ and $R^{B^{+}}$ are similar. For each $k\in B$, let
$\begin{split}B^{\prime}_{0}&=\{(0,-1,l)\mid l\in A^{\prime}\}\\ B^{\prime}_{k}&=\{(k-1,1,l)\mid l\in B^{\prime}\}\cup\{(k,-1,l)\mid l\in B^{\prime}\},\;\hbox{ for }0<k\in B,\end{split}$
and let $R^{B^{+}}$ be the smallest reflexive and symmetric relation with
$\begin{split}(k,\epsilon,l)\,&R^{B^{+}}(k,\epsilon,l^{\prime})\Leftrightarrow|l-l^{\prime}|\leq 1,\;\hbox{ for }k\in B\hbox{ and }\epsilon\in\{-1,1\},\\ (k-1,1,0)\,&R^{B^{+}}(k,-1,0),\;\hbox{ for }0<k\in B.\end{split}$
Let $B^{+}$ be the union
$\bigcup_{k\in B}B^{\prime}_{k}$
with the points $(k-1,-1,n^{\prime})\in B^{\prime}_{k-1}$ and $(k-1,1,n^{\prime})\in B^{\prime}_{k}$ identified for each $0<k\in B$. Let $R^{B^{+}}$ be the relation on $B^{+}$ that is induced from $R^{B^{+}}$ defined on $B^{\prime}_{k}$ for $k\in B$. We interpret the constant as (0,1,0)(0,-1,0).
It is easy to check that the structures $A^{+}$ and $B^{+}$ defined above are in $\mathcal{P}$$_{\bullet}$.
Let $\alpha\colon A^{+}\to A$ and $\alpha^{\prime}\colon A^{+}\to A^{\prime}$ be defined by
$\alpha(i,\epsilon,j)=i\;\hbox{ and }\;\alpha^{\prime}(i,\epsilon,j)=j.$
Similarly, let $\beta\colon B^{+}\to B$ and $\beta^{\prime}\colon B^{+}\to B^{\prime}$ be defined by
$\beta(k,\epsilon,l)=k\;\hbox{ and }\;\beta^{\prime}(k,\epsilon,l)=l$
It is easily checked that $\alpha,\alpha^{\prime},\beta,\beta^{\prime}$ are epimorphisms in $\mathcal{P}$$_{\bullet}$.
Assume now that $f\colon A\to B$ and $f^{\prime}\colon A^{\prime}\to B^{\prime}$ are epimorphisms in $\mathcal{P}$$_{\bullet}$. We need to find an epimorphism $f^{+}\colon A^{+}\to B^{+}$ in $\mathcal{P}$$_{\bullet}$ such that
[…]
By induction on $i\in A$, we define $f^{+}$ on $A_{0}^{\prime}\cup\cdots\cup A^{\prime}_{i}$. This will be done so that the value of $f^{+}$ at the last vertex of $A_{0}^{\prime}\cup\cdots\cup A^{\prime}_{i}$, which is $(i,-1,m^{\prime})$ is given by
(4.9) $f^{+}(i,-1,m^{\prime})=(f(i),\epsilon_{i},f^{\prime}(m^{\prime})),$
for some $\epsilon_{i}\in\{-1,1\}$.
We start with defining $f^{+}$ on $A^{\prime}_{0}$ by letting
$f^{+}(0,-1,j)=(0,-1,f^{\prime}(j))\;\hbox{ for }j\in A^{\prime}.$
We note that (4.9) holds with $i=0$.
Assume that $f^{+}$ is defined on $A_{0}^{\prime}\cup\cdots\cup A^{\prime}_{i-1}$ for some $0<i\in A$ and $\epsilon_{i-1}$ is determined by (4.9). We need to define $f^{+}$ on $A^{\prime}_{i}$. Let $\delta_{i}\in\{-1,0,1\}$ be such that
$f(i)=f(i-1)+\delta_{i}.$
We divide the definition of $f^{+}$ on $A^{\prime}_{i}$ into two cases. Keep in mind that we need to give the values of $f^{+}$ on triples of the form (i1,1,j)(i-1,1,j) and (i,1,j)(i,-1,j) with $j\in A^{\prime}$.
First assume $\epsilon_{i-1}\cdot\delta_{i}\geq 0$, that is, either $\delta_{i}=0$ or $\delta_{i}=\epsilon_{i-1}$ . In this case, let
$f^{+}(i-1,1,j)=(f(i-1),\epsilon_{i-1},f^{\prime}(j))$
and
$f^{+}(i,-1,j)=\begin{cases}(f(i),\epsilon_{i-1},f^{\prime}(j)),&\hbox{ if }\delta_{i}=0;\\ (f(i),-\epsilon_{i-1},f^{\prime}(j)),&\hbox{ if }\delta_{i}=\epsilon_{i-1}.\end{cases}$
Note that with these definitions (4.9) holds for $i$.
The second case is $\epsilon_{i-1}\cdot\delta_{i}<0$, that is, $\delta_{i}=-\epsilon_{i-1}$. Let
$j_{f^{\prime}}=\hbox{ the smallest }j\in A^{\prime}\hbox{ with }f^{\prime}(j_{f^{\prime}})=n^{\prime}.$
Define
$f^{+}(i-1,1,j)=\begin{cases}(f(i-1),\epsilon_{i-1},f^{\prime}(j)),&\hbox{ for }\;j_{f^{\prime}}\leq j;\\ (f(i-1),-\epsilon_{i-1},f^{\prime}(j)),&\hbox{ for }\;j\leq j_{f^{\prime}}.\end{cases}$
and
$f^{+}(i,-1,j)=(f(i),\epsilon_{i-1},f^{\prime}(j)).$
Again (4.9) holds for $i$.
Since $f^{+}$ is given by formulas, it is easy (if somewhat tedious) to check, using condition (4.9), that $f^{+}$ is an epimorphism in $\mathcal{P}$$_{\bullet}$. We leave this check to the reader. It is immediate that (4.8) holds. ∎
For the corollary below, we recall the result of Hamilton [8] that each homeomorphism of the pseudoarc has a fixed point. (In fact, each continuous map from the pseudoarc to itself has a fixed point.) The following theorem gives another strengthening of the result in [2]. It gives a homeomorphism of the pseudoarc with a dense conjugacy class with additional control on fixed points.
Corollary 4.6.
There exists $f\in{\rm Homeo}(P)$ with the following property. For each $h\in{\rm Homeo}(P)$ and $x\in P$ with $x$ being a fixed point of $h$, there exists a sequence $(g_{n})$ in ${\rm Homeo}(P)$ such that
  1. the sequence $(g_{n}fg_{n}^{-1})$ converges to $h$ in ${\rm Homeo}(P)$;
  2. $x$ is a fixed point of $g_{n}fg_{n}^{-1}$ for each $n$.
Proof..
[…]
@@ 5. A homeomorphism of the pseudoarc that is not conjugate to an automorphism of the pre-pseudoarc
This section concerns $\mathcal{P}$}—the class of finite connected linear reflexive graphs with all edge-preserving surjective homomorphisms.
@@
Let $P$ be the pseudoarc, which we identify with the natural quotient of the projective Fraïssé limit $\mathbb{P}$} of the class of finite linear graphs. There exists a homeomorphism in ${\rm Homeo}(P)$ that is not conjugate to an element of ${\rm Aut}({\mathbb{P}})$.
We note that $\mathcal{P}$} obviously fulfills condition (iv) of Proposition A.2; so it fulfills conditions (i)–(iii) in Proposition A.2 and the conclusions of Propositions A.4 and A.5. It was proved in [9] that each equivalence class of the compact equivalence relation $R^{\mathbb{P}}$ on $\mathbb{P}$} has at most two elements. This fact will be used in the proof below.
The following is a key lemma; it provides a sufficient condition for a homeomorphism not to be conjugate to any automorphism of $\mathbb{P}$}. Our argument proving the lemma hinges on the rigidity of automorphisms—they preserve the algebra of regular open sets associated with the projections from $\mathbb{P}$} onto finite linear graphs. Our example, constructed in Proposition 5.3, will satisfy the premises of the lemma.
@@
Suppose that $\sigma:P\to P$ is a homeomorphism, for which there exist a subcontinuum $C\subseteq P$ and $x\neq y\in C$ with the following properties:
  1. $\sigma{\restriction}C={\rm id}_{C}$;
  2. if $U_{1},U_{2},U_{3}$ are regular open subsets of $P$ such that
    $\begin{split}&U_{1}\cap U_{2}=U_{2}\cap U_{3}=U_{1}\cap U_{3}=\emptyset,\\ &{\rm cl}(U_{1})\cup{\rm cl}(U_{2})\cup{\rm cl}(U_{3})=P,\;\;{\rm cl}(U_{1})\cap{\rm cl}(U_{2})\cap{\rm cl}(U_{3})=\emptyset,\;\hbox{ and }\\ &{\rm cl}(U_{1})\cap C=\emptyset,\;\;x\in U_{2},\;\;y\in U_{3},\end{split}$
    then
[…]
Then $\sigma$ is not conjugate in ${\rm Homeo}(P)$ to any member of ${\rm Aut}({\mathbb{P}})$.
@@ 1. $\sigma{\restriction}C={\rm id}_{C}$;
Note that the above property of $\sigma$ is invariant under conjugation by a homeomorphism, therefore it suffices to show that representing $P$ as a quotient of the prespace $\mathbb{P}$}, we necessarily have $\sigma\notin{\rm Aut}({\mathbb{P}}$). Let ${\rm pr}:{\mathbb{P}}\to P$ be the canonical continuous surjection. Suppose that $\sigma$ is an automorphism, that is, there exists $\tilde{\sigma}\in{\rm Aut}({\mathbb{P}})$ with ${\rm pr}(\tilde{\sigma})=\sigma$, where ${\rm pr}:{\rm Aut}({\mathbb{P}})\to{\rm Homeo}(P)$ is the canonical embedding; in other words, we have
$\sigma\circ{\rm pr}={\rm pr}\circ\tilde{\sigma}$.
We set
${\mathbb{C}}={\rm pr}^{-1}(C),$
Note that $\mathbb{C}$} is a compact zero-dimensional space. We claim that
(5.1) $\tilde{\sigma}{\restriction}{\mathbb{C}}={\rm id}_{{\mathbb{C}}}.$
Indeed, since $\sigma{\restriction}C={\rm id}_{C}$, we see that $\mathbb{C}$} is invariant under $\tilde{\sigma}$ and
(5.2) $\tilde{\sigma}(u)R^{{\mathbb{P}}}u,\;\hbox{ for all }u\in{\mathbb{C}}.$
Since $\tilde{\sigma}$ is injective and each $R^{{\mathbb{P}}}$-equivalence class has at most two elements, we see that $\tilde{\sigma}\circ\tilde{\sigma}(u)=u$, for all $u\in{\mathbb{C}}$. So, if there is $u\in{\mathbb{C}}$ with $\tilde{\sigma}(u)\not=u$ for some $u\in{\mathbb{C}}$, then by taking $U$ to be a small enough clopen neighborhood of $u$ (in $\mathbb{C}$}) and $V=\tilde{\sigma}(U)$, we see that there are relatively clopen subsets $U,V$ of $\mathbb{C}$} with
(5.3) $U\cap V=\emptyset,\;\;\tilde{\sigma}(U)=V.$
and
(5.4) $U\cup V\not$$\mathbb{C}$},
so the complementer clopen set
$Z={\mathbb{C}}\setminus(U\cup V)\neq\emptyset.$
Since each $R^{{\mathbb{P}}}$-equivalence class has at most two elements, it follows from (5.2) and (5.3) that $R^{{\mathbb{P}}}(U\cup V)=U\cup V$. Since $\mathbb{C}$} is $\overline{\sigma}$-invariant, necessarily $R^{{\mathbb{P}}}(Z)=Z$. Then ${\rm pr}(U\cup V)\cap{\rm pr}(Z)=\emptyset$ contradicting that $C={\rm pr}({\mathbb{C}})$ is connected.
Fix a finite linear graph $I_{0}$ and $\varphi\in{\rm Epi}({\mathbb{P}},I_{0})$ such that $\varphi\circ{\rm pr}^{-1}(x)\cap\varphi\circ{\rm pr}^{-1}(y)=\emptyset$. Using Claim 3.9 we can find a finite linear graph $I_{1}$ and epimorphism $\psi\in{\rm Epi}({\mathbb{P}},I_{1})$ such that for some $\alpha,\alpha^{\prime}\in{\rm Epi}(I_{1},I_{0})$ we have
$\begin{array}[]{rcl}\varphi&=&\alpha\circ\psi,\\ \varphi\circ\overline{\sigma}&=&\alpha^{\prime}\circ\psi,\end{array}$
in other words,
(5.5) $\alpha\circ\psi\circ\overline{\sigma}=\alpha^{\prime}\circ\psi.$
Let $K_{1}=\psi[{\mathbb{C}}]$ (which will, roughly speaking represent $\mathbb{C}$} in $I_{1}$) we note that condition (5.1) (together with (5.5)) implies
(5.6) $\alpha{\restriction}K_{1}=\alpha^{\prime}{\restriction}K_{1}.$
Similarly, we let $K_{0}=\alpha\circ\psi[{\mathbb{C}}]$. We also let
$X_{0}=(\alpha\circ\psi)\circ{\rm pr}^{-1}(x)\;\hbox{ and }\;Y_{0}=(\alpha\circ\psi)\circ{\rm pr}^{-1}(y),$
we have $X_{0},Y_{0}\subseteq K_{0}\subseteq I_{0}$, and
(5.7) $X_{0}\cap Y_{0}=\emptyset$
by the choice of $\varphi=\alpha\circ\psi$ above. Similarly, let
$X_{1}=\psi\circ{\rm pr}^{-1}(x)\;\hbox{ and }\;Y_{1}=\psi\circ{\rm pr}^{-1}(y).$
Pick disjoint subsets $X^{\prime},Y^{\prime}\subseteq I_{0}$ with $X^{\prime}\cup Y^{\prime}=K_{0}$, $X_{0}\subseteq X^{\prime}$ and $Y_{0}\subseteq Y^{\prime}$. Set
$X^{+}=\alpha^{-1}(X^{\prime})\cap K_{1}\;\hbox{ and }\;Y^{+}=\alpha^{-1}(Y^{\prime})\cap K_{1}.$
Then we have
(5.8) $X^{+}\cup Y^{+}=K_{1},$
(5.9) $X^{+}\cap Y^{+}=\emptyset,$
(5.10) $X_{1}\subseteq X^{+},\ Y_{1}\subseteq Y^{+}.$
Set
$V_{1}=I_{1}\setminus K_{1}.$
Let
$W_{1}={\rm pr}(\psi^{-1}(V_{1})),\;\;W_{2}={\rm pr}(\psi^{-1}(X^{+})),\;\;W_{3}={\rm pr}(\psi^{-1}(Y^{+})).$
Then, by (5.8) and the definition of $V_{1}$, we have
(5.11) $W_{1}\cup W_{2}\cup W_{3}=P.$
Note furthermore that
(5.12) $W_{1}\cap W_{2}\cap W_{3}=\emptyset.$
Otherwise, there exist $w_{1}\in\psi^{-1}(V_{1})$, $w_{2}\in\psi^{-1}(X^{+})$, and $w_{3}\in\psi^{-1}(Y^{+})$ such that ${\rm pr}(w_{1})={\rm pr}(w_{2})={\rm pr}(w_{3})$, that is,
$w_{1}R^{\mathbb{P}}w_{2}R^{\mathbb{P}}w_{3}R^{\mathbb{P}}w_{1}.$
Since each $R^{\mathbb{P}}$-equivalence class has at most two elements, we see that $w_{i}=w_{j}$ for some $1\leq i<j\leq 3$, which entails that the sets $\psi^{-1}(V_{1}),\,\psi^{-1}(X^{+}),\,\psi^{-1}(Y^{+})$ are not pairwise disjoint, which contradicts pairwise disjointness of $V_{1},X^{+},Y^{+}$. The equality $K_{1}=\psi[{\rm pr}^{-1}(C)]$ together with $V_{1}\cap K_{1}=\emptyset$ imply
(5.13) $C\cap W_{1}=\emptyset.$
Also, it follows from (5.10) and the definition of $X_{1},Y_{1}$ that
(5.14) $x\notin W_{1}\cup W_{3}\;\hbox{ and }\;y\notin W_{1}\cup W_{2}.$
By Proposition A.4, we have
(5.15) ${\rm int}(W_{i})=P\setminus\bigcup_{j\not=i}W_{j}\;\hbox{ and }\;{\rm cl}\big({\rm int}(W_{i})\big)=W_{i},\;\hbox{ for }i=1,2,3.$
It follows from (5.11)–(5.15) that $U_{i}={\rm int}(W_{i})$, $i=1,2,3$, satisfy the requirements of the condition in the statement of the lemma.
@@
Now, assume that $\sigma(U_{2})\cap U_{3}\neq\emptyset$ as the other case can be dealt with the same way. Pick $z\in{\mathbb{P}}$ such that ${\rm pr}(z)\in U_{2}$ and $\sigma({\rm pr}(z))={\rm pr}(\tilde{\sigma}(z))\in U_{3}$. By the definition of $U_{2}$, we have ${\rm pr}(z)\in W_{2}\setminus(W_{1}\cup W_{3})$, so
(5.16) $\psi(z)\in X^{+},$
and, therefore,
(5.17) $\alpha(\psi(z))\in X^{\prime}.$
By a similar argument, we have $\psi(\tilde{\sigma}(z))\in Y^{+}$, and so
(5.18) $\alpha(\psi(\tilde{\sigma}(z)))\in Y^{\prime}.$
Now (5.5) and (5.6) (and (5.10)) together with (5.16) and (5.17) give
(5.19) $\alpha(\psi(\tilde{\sigma}(z)))=\alpha^{\prime}(\psi(z))=\alpha(\psi(z))\in X^{\prime}.$
But (5.18) and (5.19) contradict the disjointness of $X^{\prime}$ and $Y^{\prime}$. ∎
@@
To construct a homeomorphism that is not conjugate to any element of Aut(($\mathbb{P}$}) it remains to find one that satisfies the premises of Lemma 5.2.
@@
There exists $\sigma\in{\rm Homeo}(P)$ as in Lemma 5.2.
@@
  • define a generic sequence $J_{0},J_{1},\ldots,J_{n},\ldots$ of finite linear graphs by hand (together with the bonding maps $\pi_{i,j}$, $i\leq j<\omega$),